OpenStudy (anonymous):

What value is added to both side of the equation x^2-2x=10 in order to solve by completing the square? Please help!

OpenStudy (welshfella):

x^2-2x=10 (x - 1)^2 + a = 10 + a (let value to be added = a) if you expand (x - 1)^2 you will see what value of a you must use in order to get the left side equal to x^2 - 2x

OpenStudy (anonymous):

So it would be -2?